Firm Z is a monopolist that selss output at price $30/unit. If Firm Zs marginal cost of production is $24/unit, use the mark-up formula to find the elasticity of demand being faced by Firm Z. (Dont forget to include a minus sign in your answer!)
Answer #1

Answer :

Mark up = selling price/ Marginal cost -1

Mark up=30/24 -1=0.25

Elasticity of demand= -1/Mark up= -1/0.25=-4